### 18 Dec 2023 v2

# A Note on Minimax Robustness of Designs Against Correlated or Heteroscedastic Responses

Abstract
We present a result according to which certain functions of covariance matrices are maximized at scalar multiples of the identity matrix. This is used to show that experimental designs that are optimal under an assumption of independent, homoscedastic responses can be minimax robust, in broad classes of alternate covariance structures. In particular it can justify the common practice of disregarding possible dependence, or heteroscedasticity, at the design stage of an experiment.
keywords:
Correlation, Covariance, Induced matrix norm, Loewner ordering, Minimax, Robustness.
journal: U. Alberta preprint series
1 Introduction
Experimental designs are typically derived, or chosen, assuming that the responses will be independent and homoscedastic. As well as being simple, this is almost necessary, unless an alternative covariance structure is known. This is frequently a complicating feature of design theory – until a design is constructed and implemented there are no data which can be used to estimate an alternate structure. There is some consolation however if a proposed design, optimal under an assumption of independence and homoscedasticity, is minimax against a broad class of alternative covariance structures. By this we mean that the maximum loss in such a class is minimized by this design, which is thus robust against these alternatives. In this note we establish that any function of these alternate covariance matrices, which possesses a natural monotonicity property, is maximized within such classes at a scalar multiple of the identity matrix. This can be paraphrased by saying that the ‘least favourable’ covariance structure is that of independence and homoscedasticity. These are eventualities for which the proposed design is optimal; hence it is minimax.
2 Main result
Suppose that $\left\|\cdot\right\|_{M}$ is a matrix norm, induced by the vector norm $\left\|\cdot\right\|_{V}$, i.e.
$\left\|C\right\|_{M}=\sup_{\left\|{x}\right\|_{V}\text{ }=1}\left\|{Cx}\right\|_{V}.$ (1)
We use the subscript ‘$M$’ when referring to an arbitrary matrix norm, but adopt special notation in the following cases:
(i) For the Euclidean norm $\left\|{x}\right\|_{V}=({{x}^{T}{x}})^{1/2}$, the matrix norm $\left\|C\right\|_{M}=\left\|C\right\|_{E}$ is the spectral radius, i.e. the root of the maximum eigenvalue of $C^{T}C$. This is the maximum eigenvalue of $C$ if $C$ is a covariance matrix, i.e. is symmetric and positive semidefinite.
(ii) For the sup norm $\left\|{x}\right\|_{V}=\max_{i}\left|x_{i}\right|$, the matrix norm $\left\|C\right\|_{M}=\left\|C\right\|_{\infty}$ is $\max_{i}\sum_{j}\left|c_{ij}\right|$, the maximum absolute row sum.
(iii) For the 1-norm $\left\|{x}\right\|_{V}=\sum_{i}\left|x_{i}\right|$, the matrix norm $\left\|C\right\|_{M}=\left\|C\right\|_{1}$ is $\max_{j}\sum_{i}\left|c_{ij}\right|$, the maximum absolute column sum. For symmetric matrices, $\left\|C\right\|_{1}=\left\|C\right\|_{\infty}$.
See Todd (1977, Chapter 3) for verifications of (i) – (iii).
We require the following properties of induced norms:
P1) $\left\|I\right\|_{M}=1$.
P2) For covariance matrices $C$, $\left\|C\right\|_{M}\geq\left\|C\right\|_{E}$.
Property P1) is immediate from (1). For P2), suppose that $\left\|C\right\|_{M}$ is induced by $\left\|\cdot\right\|_{V}$, and that $\lambda_{0}$ is the maximum eigenvalue of $C$, with eigenvector $x_{0}$ normalized so that $\left\|x_{0}\right\|_{V}=1$. Then
$\left\|C\right\|_{E}=\lambda_{0}=\left\|\lambda_{0}x_{0}\right\|_{V}=\left\|Cx_{0}\right\|_{V}\leq\sup_{\left\|x\right\|_{V}=1}\left\|Cx\right\|_{V}=\left\|C\right\|_{M}.$
Now suppose that the loss function in a statistical problem is $\mathcal{L}\left(C\right)$, where $C$ is an $N\times N$ covariance matrix and $\mathcal{L}\left(\mathbf{\cdot}\right)$ is non-decreasing in the Loewner ordering:
$A\preceq B\Rightarrow\mathcal{L}\left(A\right)\leq\mathcal{L}\left(B\right).$ (2)
Here $A\preceq B$ means that $B-A\succeq 0$, i.e. is positive semidefinite.
Lemma 1.
For any positive constant $\eta^{2}$, covariance matrix $C$ and norm $\left\|C\right\|_{M}$, define
$\mathcal{C}_{M}=\{C\mid C\succeq 0\text{ and }\left\|C\right\|_{M}\leq\eta^{2}\}.$
For the norm $\left\|\mathbf{\cdot}\right\|_{E}$ an equivalent definition is
$\mathcal{C}_{E}=\{C\mid 0\preceq C\preceq\eta^{2}I_{N}\}.$ (3)
Then:
(i) In any such class $\mathcal{C}_{M}$, $\max_{\mathcal{C}_{M}}\mathcal{L}\left(C\right)=\mathcal{L}\left(\eta^{2}I_{N}\right)$.
(ii) If $\mathcal{C^{\prime}}$ $\mathcal{\subseteq C}_{M}$ and $\eta^{2}I_{N}\in\mathcal{C^{\prime}}$, then $\max_{\mathcal{C^{\prime}}}\mathcal{L}\left(C\right)=\mathcal{L}\left(\eta^{2}I_{N}\right)$.
Proof.
We first establish the equivalence of (3). If the condition there holds, then by Weyl’s Monotonicity Theorem (Bhatia 1997, p. 63), all eigenvalues of $C$ are dominated by $\eta^{2}$, hence $\left\|C\right\|_{E}\leq\eta^{2}$. Conversely, if $\left\|C\right\|_{E}\leq\eta^{2}$ then all eigenvalues of $C$ are dominated by $\eta^{2}$, hence all those of $\eta^{2}I_{N}-C$ are non-negative and so the condition in (3) holds.
[…]
Remark 1.
An interpretation of Lemma 1 is as follows. Suppose that one has derived a technique under an assumption of uncorrelated, homoscedastic errors, i.e. a covariance matrix $\sigma^{2}I_{N}$, which is optimal in the sense of minimizing $\mathcal{L}$, for any $\sigma^{2}>0$. Now suppose one is concerned that the covariance matrix might instead be a member $C$ of $\mathcal{C}_{M}$, and that $\mathcal{L}$ is monotonic in the sense described above. Then the technique minimizes $\max_{C_{M}}\mathcal{L}\left(C\right)=\mathcal{L}\left(\eta^{2}I_{N}\right)$, i.e. is minimax in $\mathcal{C}_{M}$.
Remark 2.
In Remark 1 we implicitly assume that $\eta^{2}\geq\sigma^{2}$, else $\mathcal{C}_{M}$ does not contain $\sigma^{2}I_{N}$.  An argument for taking $\eta^{2}>\sigma^{2}$ arises if one assumes homoscedasticity and writes $C=\sigma^{2}P$, where $P$ is a correlation matrix. Then in $\mathcal{C}_{1}$, $\eta^{2}\geq\left\|C\right\|_{1}$ $=\sigma^{2}\left\|P\right\|_{1}\geq\sigma^{2}$, with the final inequality being an equality iff $P=I_{N}$. Thus take $\eta^{2}>\sigma^{2}$. Then an intuitive explanation of the lemma is that in determining a least favourable covariance structure, one can alter the correlations in some manner which increases $\left\|C\right\|_{M}$, or one can merely increase the variances. The answer is that one should always just increase the variances.
Remark 3.
A version of Lemma 1 was used by Wiens and Zhou (2008) in a maximization problem related to the planning of field experiments. It was rediscovered by Welsh and Wiens (2013) in a study of model-based sampling procedures. This note seems to be the first systematic study of the design implications of the lemma.
3 Applications
3.1 Experimental design in the linear model
Consider the linear model $y=X\theta+\varepsilon$. Suppose that the random errors $\varepsilon$ have covariance matrix $C\in\mathcal{C}_{M}$. If $C$ is known then the ‘Best Linear Unbiased Estimate’ is $\hat{\theta}_{\text{{blue}}}=$ $\left(X^{T}C^{-1}X\right)^{-1}X^{T}C^{-1}y$ and there is an extensive design literature - see Dette et al. (2015) for a review. In the more common case that the covariances are only vaguely known, or perhaps only suspected, it is more usual to use the ordinary least squares estimate $\hat{\theta}_{\text{{ols}}}$, design as though the errors are uncorrelated, and hope for the best. An implication of the results of this section is that, in a minimax sense, that approach can be quite sensible.
In the classical ‘alphabetic’ design problems, one seeks to minimize a function $\Phi$ of the covariance matrix of the regression estimates. Let $\xi_{0}$ be the minimizing design, under the possibly erroneous assumption of uncorrelated, homoscedastic errors. Assume that under $\xi_{0}$the moment matrix $X^{T}X$ is non-singular. Then the covariance matrix of $\hat{\theta}_{\text{{ols}}}$ is
$\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C)=\left(X^{T}X\right)^{-1}X^{T}CX\left(X^{T}X\right)^{-1}.$ (4)
Suppose that $0\preceq C_{1}\preceq C_{2}$, so that $C_{2}-C_{1}=A^{T}A$ for some $A$. Then
$\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{2})-\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{1})=B^{T}B\succeq 0,$
for $B$ $=AX\left(X^{T}X\right)^{-1}$, hence $\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{1})\preceq\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{2}).$ It follows that if $\Phi$ is non-decreasing in the Loewner ordering, then $\mathcal{L}(C)=\Phi\{\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C)\}$ is also non-decreasing and the conclusions of the lemma hold. Then as at Remark 1, $\xi_{0}$ is a minimax design – it minimizes the maximum loss as the covariance structure varies over $\mathcal{C}_{M}$.
Again by Weyl’s Monotonicity Theorem, if $0\preceq\Sigma_{1}\preceq\Sigma_{2}$ then the $i$th largest eigenvalue $\lambda_{i}$ of $\Sigma_{2}$ dominates that of $\Sigma_{1}$, for all $i$. It follows that $\Phi$ is non-decreasing in the Loewner ordering if:
(i) $\Phi\left(\Sigma\right)=tr\left(\Sigma\right)=\sum_{i}\lambda_{i}\left(\Sigma\right)$, corresponding to ‘A-optimality’;
(ii) $\Phi\left(\Sigma\right)=det\left(\Sigma\right)=\prod_{i}\lambda_{i}\left(\Sigma\right)$, corresponding to ‘D-optimality’;
(iii) $\Phi\left(\Sigma\right)=\max_{i}\lambda_{i}\left(\Sigma\right)$, corresponding to ‘E-optimality’;
(iv) $\Phi\left(\Sigma\right)=tr\left(L\Sigma\right)$ for $L\succeq 0$, corresponding to ‘L-optimality’ and including ‘I-optimality’ – minimizing the integrated variance of the predictions;
thus the designs optimal under any of these criteria are minimax.
3.1.1 ma(1) (moving average of order 1) errors.
As a particular case, assume first that the random errors are homoscedastic but are possibly serially correlated, following an ma(1) model with corr$(\varepsilon_{i},\varepsilon_{j})=\rho I\left(\left|i-j\right|=1\right)$ and with $\left|\rho\right|\leq\rho_{\max}$. Then under this structure $C$ varies over the subclass $\mathcal{C}^{\prime}$ of $\mathcal{C}_{\infty}$ defined by $c_{ij}=0$ if $\left|i-j\right|>1$ and $\left\|\boldsymbol{C}\right\|_{\infty}\leq\sigma^{2}\left(1+2\rho_{\max}\right)$, which we define to be $\eta^{2}$. Since $\eta^{2}I_{N}\in\mathcal{C}^{\prime}$, part (ii) of the lemma applies and yields that $\xi_{0}$ is a minimax design in $\mathcal{C}^{\prime}$ and with respect to any of the alphabetic criteria above. If the errors are instead heteroscedastic, then $\sigma^{2}$ is replaced by the maximum of the assumed variances.
3.1.2 ar(1) (autoregressive of order 1) errors.
It is known that the eigenvalues of an ar(1) autocorrelation matrix with autocorrelation parameter $\rho$ are bounded, and that the maximum eigenvalue $\lambda_{N}\left(\rho\right)$ has $\lambda_{N}^{\ast}=$ $\max_{\rho}\lambda_{N}\left(\rho\right)>\lambda_{N}\left(0\right)=1$. See for instance Trench (1999, p. 182). Then, again under homoscedasticity, the covariance matrix $C$ has $\left\|C\right\|_{E}\leq\sigma^{2}\lambda_{N}^{\ast}$, and a design optimal when $\rho=0$ is minimax in the subclass of $\mathcal{C}_{E}$ defined by the autocorrelation structure and $\eta^{2}=\sigma^{2}\lambda_{N}^{\ast}$.
3.2 Designs robust against model misspecifications
Working in finite design spaces $\chi$ and with $p$-dimensional regressors $f\left(x\right)$, Wiens (2018) derived minimax designs for possibly misspecified regression models
$Y\left(x\right)=f^{T}\left(x\right)\theta+\psi\left(x\right)+\varepsilon,$ (5)
with the unknown contaminant $\psi$ ranging over a class $\Psi$ and satisfying, for identifiability of $\theta$, the orthogonality condition
$\sum_{x\in\chi}f\left(x\right)\psi\left(x\right)=0_{p\times 1}.$ (6)
For designs $\xi$ placing mass $\xi_{i}$ on $x_{i}\in\chi$, he took $\hat{\theta}=\hat{\theta}_{\text{{ols}}}$,
$\mathcal{I}\left(\psi,\xi\right)$ $=$ $\sum_{x\in\chi}E[f^{T}\left(x\right)\hat{\theta}-E\{Y\left(x\right)\}]^{2},$ (7a)
$\mathcal{D}\left(\psi,\xi\right)$ $=$ $[\det E\{(\hat{\theta}-\theta)(\hat{\theta}-\theta)^{T}\}]^{1/p},$ (7b)
and found designs minimizing the maximum, over $\psi$, of these loss functions. The random errors $\varepsilon_{i}$ were assumed to be i.i.d.; now suppose that they instead have covariance matrix $C\in\mathcal{C}_{M}$.
Consider first (7a). Using (6), and with $d_{\psi}=\{E(\hat{\theta})-\theta\}$, which does not depend on the covariance structure, this decomposes as
$\mathcal{I}\left(\psi,\xi,C\right)=\sum_{x\in\chi}f^{T}\left(x\right){\textsc{cov}}(\hat{\theta}\mid C)f\left(x\right)+\sum_{x\in\chi}\left\{f^{T}\left(x\right)d_{\psi}d_{\psi}^{T}f\left(x\right)+\psi^{2}\left(x\right)\right\}.$ (8)
The first sum above does not depend on $\psi$; the second depends on $\psi$ but not on the covariance structure. Then an extended minimax problem is to find designs $\xi$ minimizing
$\max_{\psi,C}\mathcal{I}\left(\psi,\xi,C\right)=\max_{C\in\mathcal{C}_{M}}\sum_{x\in\chi}f^{T}\left(x\right)\text{{cov}}(\hat{\theta}\mid C)f\left(x\right)+\max_{\psi\in\Psi}\sum_{x\in\chi}\left\{f{}^{T}\left(x\right)d_{\psi}d_{\psi}^{T}f\left(x\right)+\psi^{2}\left(x\right)\right\}.$
As in §3.1, and taking $L=\sum_{x\in\chi}f\left(x\right)f^{T}\left(x\right)$ in (iv) of that section, for $C\in\mathcal{C}_{M}$ the first sum is maximized by a multiple $\eta^{2}$ of the identity matrix, and then the remainder of the minimax problem is that which was solved in Wiens (2018). The minimax designs, termed ‘I-robust’ designs, obtained there do not depend on the value of $\eta^{2}$, and so enjoy the extended property of minimizing $\max_{\psi,C}\mathcal{I}\left(\psi,\xi,C\right)$ for $C\in\mathcal{C}_{M}$.
Now consider (7b). The analogue of (8) is $\mathcal{D}\left(\psi,\xi,C\right)=[\det\{\text{{cov}}(\hat{\theta}\mid C)+d_{\psi}d_{\psi}^{T}\}]^{1/p}.$ Since cov$(\hat{\theta}\mid C)+d_{\psi}d_{\psi}^{T}$, hence its determinant, is non-decreasing in the Loewner ordering, $\mathcal{D}\left(\psi,\xi,C\right)$ is maximized, for $C\in\mathcal{C}_{M}$, by a multiple of the identity matrix. The rest of the argument is identical to that in the preceding paragraph, and so the ‘D-robust’ designs obtained in Wiens (2018) also minimize $\max_{\psi,C}\mathcal{D}\left(\psi,\xi,C\right)$ for $C\in\mathcal{C}_{M}$.
Remark 4.
Results in the same vein as those above have been obtained in cases which do not seem to be covered by Lemma 1. For instance Wiens and Zhou (1996) sought minimax designs for the misspecification model (5), under conditions on the spectral density of the error process. They state that “… a design which is asymptotically (minimax) optimal for uncorrelated errors retains its optimality under autocorrelation if the design points are a random sample, or a random permutation, of points …”, with details in their Theorems 2.4 and 2.5.
3.3 Designs for nonlinear regression models
In the nonlinear regression model $y=f\left(x;\theta\right)+\varepsilon$, the goal of a design is typically the minimization of some function of the covariance matrix of $\hat{\theta}$, after a linearization of the model. When the errors have covariance $C$ the target of this minimization continues to be given by the right hand side of (4), but with the matrix $X$ replaced by the gradient $F\left(\theta\right)=\{\partial f\left(x_{i};\theta\right)/\partial\theta_{j}\}_{i,j}$, and with $F\left(\theta\right)$ evaluated at an initial parameter estimate, or at a previous estimate in an interative estimation scheme. Then the results of §3.1 continue to hold, and a design optimal for the loss functions given there, under independence and homoscedasticity, is minimax. A caveat, however, is that optimal design theory for nonlinear models is much less well-developed, and more model specific, than that for linear models. See Bates and Watts (1988) and Hamilton and Watts (1985) for discussions.
Acknowledgement
This work was carried out with the support of the Natural Sciences and Engineering Council of Canada. It has benefited from the helpful comments of two anonymous referees.